# Hidden long-range correlations in the ion distribution at the graphite / [bmim][NTf2] electrified interface

### v1

@@ I Introduction
The long range of the Coulomb charge-charge interactions poses strict conditions to the structure and dynamics of simple electrolytes in contact with an electrode. In particular, the width of the spatial charge separation at the interface is limited by screening to a few characteristic lengths, i.e., multiples of a screening length $\lambda$, whose value depends primarily on the thermodynamic state of the electrolyte. Beyond this (usually microscopic) length, the electrolyte appears to be largely bulk-like.
Given this generally accepted picture, it came as a surprise the revelation by optical measurements of characteristic shifts in the frequency of characteristic lines, affecting the spectroscopic properties of IL electrolytes far from the interface.expLangmuir
@@
Since the strict confinement of the interfacial charge perturbation is supported by very robust arguments and is verified by a large number of experimental, theoretical and computational studies, we look for the responsible of this long range ordering at the second simpler and most likely suspect, represented by the organisation of dipoles by long range correlation effects. As the testbed for this analysis we have chosen the interface between a graphite electrode and a paradigmatic IL, consisting of the 1-butyl-3-methylimidazolium cation ([bmim]+) and bis(trifluoromethylsulfonyl)-imide anion ([NTf2]-), i.e., [bmim][NTf2]. This compound is mildly hydrophobic but retains a not-completely-negligible water concentration of up to about $25$ molar %, which however corresponds to only less that $1.5$ wgt% due to the mass disparity of water and [bmim][NTf2]. We chose this system for our investigation, despite its complexity, because of its interest as a solvent electrolyteelectrochem and relatively abundance of experimental data on its physical chemistry properties. Moreover, the same effect might not be observable in simpler (either real or idealised) models because size and shape of the electrolyte ions might play a role in the propagation of correlation effects far from the interface.
@@
The first target of this investigation is to define a population of instantaneous dipoles made by the IL ions in the least arbitrary way. Achieving this goal will allow us to quantify correlations among dipoles as a function of the electron density on the electrode, of temperature and of concentration of water contamination, due to the hygroscopic character of [bmim][NTf2]. The MD of the interface, that represents the required background of this investigation, provides a wealth of additional information on the electrostatic properties of the interface, including an estimation of the differential capacitance of the interface, and its dependence on temperature, pressure, surface charge density on the electrode, and concentration of contaminant water. Since the discussion of all these aspects could obscure the major novelty of the present study, we defer the full presentation of our investigation to a later and more comprehensive paper, devoting the present document to a concise report of the main result, represented by the presence of dipole correlations far from the interface, where the average charge density and electric field vanish.
@@ II Model and Method
The simulated system is a planar capacitor consisting of two parallel graphite plates orthogonal to the $\hat{x}$ axis, delimiting a relatively thick ($\sim 11$ nm) [bmim][NTf2] layer, see Fig. 1. Both graphite and the IL are described at the atomistic level. The potential energy surface of the system is represented through an empirical force field of the Amber form,amber parametrised at first using the Gromos model in its 54a754a7 version for the [bmim][NTf2] component, gromos and the SPC model for water.spc In practice, the topology files for the [bmim]+ and [NTf2]- ions have been obtained using the automated topology building utility made available by the ATB website.atb At first, all the charges on the graphite carbon atoms are set to zero.
@@
Test simulations of homogeneous (3D periodic boundary conditions, pbc) [bmim][NTf2] systems have shown that the self-diffusion constant of both [bmim]+ and [NTf2]- as computed by the Einstein relation are severely underestimated. This is a know problem of unpolarisable IL force fields,diffu such as Gromos 54a7, and, in the present case the problem might be exacerbated by a non-negligible over-estimation of the homogeneous IL density. The diffusivity problem sometimes is solved by scaling all atomic charges by a 0.80.90.8-0.9 factor,08 ; 09 justified in terms of atomic polarizability or partial charge transfer among cations and anions (which can also be seen as a form of electronic polarizability). These effects, however, are likely to be different in the bulk and at an interface. Moreover, any ion that (hypothetically) leaves the condensed phase needs to have an integer charge. Lacking a suitable model to include the charge dependence on configuration and thermodynamic state in a consistent and computationally convenient form, it has been decided to keep unchanged the formal charge +1e+1e, 1e-1e of cations and anions, respectively. This decision implies accepting a general slowing down of the ion dynamics with respect to the experimental system. The density over-estimation, instead, has been addressed by an empirical tuning of the Lennard-Jones (LJ) pair potential, hoping to improve at the same time the diffusivity aspect. To this aim, all Lennard-Jones radii $\sigma_{ij}$ of [bmim][NTf2] atoms have been expanded by a constant factor of $1.03$ (+3%), recovering the experimental density of pure and homogeneous [bmim][NTf2] at $T=300$ K. The homogeneous liquid properties of [bmim][NTf2] estimated by this model are reported and briefly discussed in Sec. III. The aromatic carbon description of graphite has been left unchanged at the Gromos 54a7 level, since it has little effect on the IL density and dynamics even in its proximity besides confining the IL to the inside of the model capacitor.
Figure 3: [bmim][NTf2] neutral ion pair. The picture shows a generic configuration from a MD trajectory at $T=300$ K.
@@
All planar capacitor samples are enclosed in an orthorhombic box, with periodic boundary conditions applied. More precisely, at first samples are periodically replicated in 2D along the $(yz)$ plane. Then, despite the inherent 2D character of the capacitor, the electrostatic interactions have been computed using 3D Ewald sums upon replicating the capacitor also along the $\hat{x}$ direction, because 2D Ewald sum routines are generally considered to be less efficient than their 3D counterparts. No explicit test, however, has been carried out to quantify or even verify this statement for the present case. The thickness of the graphite slab ($\sim 30$ Å ) limits the spurious interactions of the two interfaces that, because of pbc along $x$ find themselves weakly interacting through each graphite block.
The artificial periodicity along $x$, however, has a far more significant effect, that needs to be accounted for. An isolated, single planar capacitor of charge $Q=\sigma A$, where $\sigma$ is the surface charge density on the positive electrode and $A$ is the electrode area, is globally neutral, and the electric field on its outside vanishes. Hence, application of Gauss’ theorem immediately gives that the electric field strength on its inside is a constant equal to $4\pi\sigma$. For an infinitely periodic sequence of capacitors, however, the analysis is not so simple, because there is no simple way to locate a surface of vanishing electric field (${\bf E}=0$) from which to start applying Gauss’ theorem to compute ${\bf E}(x)$ anywhere else in space. For this reason, the electric field inside the capacitor, although still constant and directed along $x$, is not equal to $4\pi\sigma$.chris The precise value of ${\bf E}(x)$ can be computed analytically based on the Ewald sum formalism, but it depends on boundary condition parameters set in the simulation code, whose value in not easily determined. To avoid this ambiguity, the electric field $^{\bf E_{0}}=E_{0}(x)\hat{x}$ along the $x$ coordinate within the capacitor and within the graphite slab has been computed using Gromacs considering an empty capacitor whose charged plates (represented by the charge graphite layers) are located at the distance used in the actual simulations with the ionic liquid. This empty capacitor curve $E_{0}(x)$ has been used throughout the paper to compute the position-dependent electric field ${\bf E}(x)$ and electrostatic potential ${\bf\phi}(x)$ through the capacitor.
@@
All simulated systems consist of the same number of cations and anions. The electrolyte slab, therefore, is globally neutral. Since also the whole sample, which is 3D periodic, needs to be neutral, the two interfaces delimiting the electrolyte may only carry zero or opposite surface charge densities $\pm\sigma$. At first, the charge on each graphite C atom is identically zero. Then, any non-vanishing charge density on the graphite surfaces will be represented by atomic charges $q_{grf}$ of equal modulus and opposite sign on the two surfaces. Since graphite is a conductor, the static $\pm\sigma$ charges on the two sides of the graphite, in reality, would not be stable with respect to charge recombination. However, the electrical conductivity perpendicular to the graphite planes is about $10^{3}$ times lower than along the planes. Therefore, the simple charging model is justified by neglecting the electrical conductivity across the graphite slab, and assuming that the two outermost C-planes are connected to the two poles of a battery.
All simulations have been carried out using the Gromacsgromacs simulation package, version 2019.3. The analysis of trajectories has been made using home-made programs. These have been used to compute number and charge density profiles along the direction normal to the graphite planes, ion-ion radial distribution functions, self-diffusion coefficient based on the Einstein equation, orientation correlations among the ions and with respect to the planar IL/graphite interfaces. The electrostatic potential along the same normal direction has been computed from the charge density profile by integrating the electric field determined through Gauss theorem.
Figure 4: Snapshot of the ionic liquid part of Fig. 1 coarse grained through Eq. 1
@@
whose definition corresponds to the centre of mass of a system of charges, whose mass is equal to the absolute value of their charge. This definition is not as generally accepted nor extensively used as that of the centre of mass. In this context, however, it provides a simple way to coarse grain the structure of [bmim][NTf2], which is fairly complex and difficult to visualise and interpret in its atomistic detail. For instance, the configuration of cations and anions corresponding to the snapshot in Fig. 1 is shown in Fig. 4, highlighting the drastic simplification of the IL structure achieved by the coarse graining expressed by Eq. 1
@@ III.1 Properties of the homogeneous samples
Since the experimental density was fitted by increasing the LJ radii, the model reproduces the experimental density. Its temperature dependence is slightly underestimated. At the $20$ molar % concentration, water appears to be soluble in the re-parametrised [bmim][NTf2] model, in agreement with experiments.
Figure 5: Number (blue line) and charge (red line) density profiles across the uncharged ($\sigma=0$) capacitor.
@@
Figure 6: Number (blue line) and charge (red line) density profiles across the capacitor with $\sigma=2$ e/nm2 surface charge density. Notice the change of scale with respect to Fig. 5.
@@ III.2 Structural, dynamic and electrostatic properties of the graphite / [bmim][NTf2] interface
The reference frame has been oriented in such a way that the interface is orthogonal to the $x$ axis. For this reason, local properties such as density and electrostatic potential profiles are plotted as a function of $x$. To simulate charged capacitors, small charges of up to a few $10^{-2}$ e (where e is the H+ charge) have been equally distributed among all the atoms of the two most external layers of graphite, qualitatively reflecting the picture provided by ab-initio (density functional theory) computations, see charging, . The charges are of equal magnitude and opposite sign on the two charged layers. Since also the IL slab is globally neutral, the whole sample is neutral.
Figure 7: Electric field and electrostatic potential as a function of the $x$ coordinate across the capacitor for two different values of the surface charge $\sigma$.
@@ III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
The results of the previous subsections show that the average electric field vanishes away from the interfaces. The presence of Stark effects sufficiently strong to be detected by spectroscopy, however, does not require a non-vanishing static electric field. A detectable Stark effect is in fact due to a change in the average of the square modulus $E^{2}$ of the electric field acting on each ion. Such a quantity does not vanish even in bulk [bmim][NTf2], and changes in its value can result from subtle correlations compatible with the strict screening sum rules that exclude net charge separation and non-vanishing average electric field away from the interface. Thus, excluding the violation of these rules, the present investigation focuses on possible correlations among electrostatic dipoles.
@@
Such an analysis can follow a variety of leads, but two aspects need to be retained. First, one needs to define dipoles inside an ionic fluid, which is certainly a challenging undertaking since charges in [bmim][NTf2] are mobile. Second, each dipole needs to have a compact geometry in such a way that its position in space can be specified with good precision, allowing the meaningful computation of a local density of dipoles everywhere through the system as well as the definition of distance dependent correlation functions. The subdivision of the system charges into dipoles cannot but also does not need to be permanent, but it will evolve with the changing configuration of ions in the ionic liquid. However, to lend credibility to the physical relevance of these dipoles, their lifetime should be long. Moreover, the probability distribution for the modulus $R_{d}$ over the dipoles population should be relatively narrow, and correspond to a charge separation comparable to the average distance of ions as given by a ion-ion pair distribution function.
@@
With these considerations in mind, we set out to partition the moving charges of the system into a population of dipoles. The dipole moment of a set of charges is unambiguously defined provided the aggregation of charge is overall neutral. Therefore, we focus our attention on the dipole of neutral ion pairs, consisting of the association of one cation and one anion. The difficulty with this approach is that the way to partition an assembly of ions into a collection of dipoles by joining a [bmim]+ cation and an [NTf2]- anion is not unique. The most intuitive approach could be to join each ion with its nearest neighbour of the opposite sign. However, the property of ion A+ being the nearest neighbour of B- is not symmetric, since the nearest cation of B- might be C+, different from A+. Simple algorithms like starting from a given ion, joining it to the closest counter-ion not already engaged into a dipole do not solve the problem. At first, this approach provides a series of tightly bound dipoles, satisfying the requirements listed above. Soon, however, going down the series of cations, the inherent frustration in the process of pairing neighbouring ions becomes manifest since the algorithm increasingly joins pairs separated by long distances, making meaningless the location of the dipole in space.
@@
The approach that is proposed and tested here is to define such a unique partition of the ion population into dipoles by requesting that their total self-energy is a minimum among all possible subdivision of all cations and anions into neutral pairs.
Figure 8: Snapshots of the dipole configurations at the two extreme temperatures of the simulated annealing. The $x$ axis is orthogonal to the two IL - graphite interfaces, with graphite not shown in the picture. Dark gray arrows point to the left, red arrows point to the right. $T$ is a fictitious temperature introduced to drive the MC optimisation of the ions’ pairing into dipoles, not to be confused with the physical temperature (T=300 K) of the graphite - liquid [bmim][NTf2] simulation. The units of the fictitious temperature are discussed in the text.
Let us first consider the case of neutral graphite surfaces. The case of charged interfaces will require some slight modifications. The self energy of a dipole made of charges (q,q)(q,-q) separated by a distance $s$ usually is given as $u_{s}=-q^{2}/s$. In the present case, in which cations and anions consists of several atoms joined into a complex ion, the self energy of the dipole made of cation $\alpha$ and anion $\beta$ is defined as:
@@
where $i$ and $j$ label atoms (of charges $q_{i}$ and $q_{j}$, respectively) belonging to cation $\alpha$ and anion $\beta$, respectively. To carry out the unbiased subdivision of any given configuration of ions into a configuration of dipoles, a procedure has been devised which:
  • starts from the pairing of each cation with its nearest anion which is not already engaged in another pair;

### v2

@@ I Introduction
The long range of the Coulomb charge-charge interactions poses strict conditions on the structure and dynamics of the electrostatic double layer that forms at the interface between an electrolyte and an electrode. In particular, the width of the spatial charge separation at the interface is limited by screening to a few characteristic lengths, i.e., multiples of a screening length $\lambda$, whose value depends primarily on the thermodynamic state of the electrolyte. Beyond this (usually microscopic) length, the electrolyte appears to be largely bulk-like.
Given this generally accepted picture, it came as a surprise the revelation by optical measurements of characteristic changes in the Raman response of imidazolium-based ionic liquid enclosed between the plates of a static capacitor, driven by the applied electric field and affecting the spectroscopic properties of the IL electrolytes far from the interfaces.CARS ; expLangmuir To our knowledge, up to now, effects of this type have been revealed on electrolyte samples made of:expLangmuir 1-butyl-3-methyl-imidazolium bis(trifluoromethylsulfonyl)imide ([bmim][NTf2]); 1-butyl-3-methyl-imidazolium hexafluorophosphate ([bmim][PF6]); 1-ethyl-3-methyl-imidazolium bis(trifluoromethylsulfonyl)imide ([emim][NTf2]) as well as 1-alkyl-3-methylimidazolium hexafluorophosphates ([Cnmim][NTf2], $n=4$, 66, 88).CARS
@@
Since the strict confinement of the interfacial charge perturbation is supported by very robust arguments and is verified by a large number of experimental, theoretical and computational studies, we look for the responsible of this long range ordering at the second simpler and most likely suspect, represented by the of ions into neutral dipolar aggregates, which could display long range correlation effects. As the testbed for this analysis we have chosen the interface between a graphite electrode and a paradigmatic IL, consisting of the 1-butyl-3-methylimidazolium cation ([bmim]+) and bis(trifluoromethylsulfonyl)-imide anion ([NTf2]-), i.e., [bmim][NTf2]. This compound is mildly hydrophobic but retains a not-completely-negligible water concentration of up to about $25$ molar %, which however corresponds to only less that $1.5$ wgt% due to the mass disparity of water and [bmim][NTf2]. Samples of dry and water-contaminated [bmim][NTf2] samples with and without confining graphite electrodes have been simulated by molecular dynamics (MD) using an atomistic, semi-empirical force field model. We chose [bmim][NTf2] for our investigation because, despite its complexity, it has received a considerable interest as a solvent electrolyteelectrochem and because relatively abundant experimental and computational literature is available on its physical-chemistry properties. Moreover, the same effect might not be observable in simpler (either realistic or idealised) models because size and shape of the electrolyte ions might play a role in the propagation of correlation effects far from the interface. On the other hand, graphite has been selected because it is indeed used to make electrodes, and especially because its interaction with organic species are covered by the same type of interatomic force fields used for the ions themselves, unlike metal electrodes, whose typical empirical potential energy models are not easy to combines with those of ILs.
@@
The main contribution of the present study is the development and calibration of a novel approach to analyse the MD trajectories, which, for each configuration, defines a population of instantaneous dipoles made from IL ions in the least arbitrary way. Achieving this goal will allow us to quantify correlations among dipoles as a function of the electron density on the electrode, of temperature and of concentration of water contamination, whose presence is due to the hygroscopic character of [bmim][NTf2]. The MD of the interface, that represents the required background of this investigation, provides a wealth of additional information on the electrostatic properties of the interface, including an estimation of the differential capacitance of the interface, and its dependence on temperature, pressure, surface charge density on the electrode, and concentration of contaminant water. Since the discussion of all these aspects could obscure the major novelty of the present study, we defer the full presentation of our investigation to a later and more comprehensive paper, devoting the present document to a concise report of the main result, represented by the presence of dipole correlations far from the interface, where the average charge density and electric field vanish.
@@ II Model and Method
The subject of the major portion of the simulations is a planar capacitor consisting of two parallel graphite plates orthogonal to the $\hat{x}$ axis, delimiting a relatively thick ($\sim 11$ nm) [bmim][NTf2] layer, see Fig. 1. The thickness of the [bmim][NTf2] slab is such to overlap, at least to a limited extent, with the range (1010010-100 nm) of the long range correlations found in similar experimental systems. Both graphite and the IL are described at the atomistic level. The potential energy surface of the system is represented through an atomistic, empirical force field of the Amber form,amber parametrised at first using the Gromos model in its 54a754a7 version for the [bmim][NTf2] component, gromos and the SPC model for water.spc In practice, the topology files for the [bmim]+ and [NTf2]- ions have been obtained using the automated topology building utility made available by the ATB website.atb At the beginning, all the charges on the graphite carbon atoms are set to zero, to represent uncharged electrodes. In the case of electrodes of non-vanishing surface density $\sigma$, the atomic-like charges of graphite are distributed in contributions $\pm\delta q$ of equal size and opposite sign on the outermost layers of the graphite block, as detailed in the following Sec. III.2.
@@
Test simulations of homogeneous (3D periodic boundary conditions, pbc) [bmim][NTf2] systems have shown that the self-diffusion constant of both [bmim]+ and [NTf2]- as computed by the Einstein relation are severely underestimated. This is a know problem of unpolarisable IL force fields,diffu such as Gromos 54a7, and, in the present case the problem might be exacerbated by a non-negligible over-estimation of the homogeneous IL density. The diffusivity problem sometimes is solved by scaling all atomic charges by a 0.80.90.8-0.9 factor,08 ; 09 justified in terms of atomic polarizability or partial charge transfer among cations and anions (which can also be seen as a form of electronic polarizability). These effects, however, are likely to be different in the bulk and at an interface. Moreover, any ion that (hypothetically) leaves the condensed phase needs to have an integer charge. Lacking a suitable model to include the charge dependence on the instantaneous configuration and on the thermodynamic state in a consistent and computationally convenient form, it has been decided to keep unchanged the formal charge +1e+1e, 1e-1e of cations and anions, respectively. This decision implies accepting a general slowing down of the ion dynamics with respect to the experimental system. The density over-estimation, instead, has been addressed by an empirical tuning of the Lennard-Jones (LJ) pair potential, hoping to improve at the same time the diffusivity aspect. To this aim, all Lennard-Jones radii $\sigma_{ij}$ of [bmim][NTf2] atoms have been expanded by a constant factor of $1.03$ (+3%), recovering the experimental density of pure and homogeneous [bmim][NTf2] at $T=300$ K. The homogeneous liquid properties of [bmim][NTf2] estimated by this model are reported and briefly discussed in Sec. III. The aromatic carbon description of graphite has been left unchanged at the Gromos 54a7 level, since it has little effect on the IL density and dynamics even in its proximity besides confining the IL to the inside of the model capacitor.
Figure 3: [bmim][NTf2] neutral ion pair. The color coding is the same as in Fig. 1 The picture shows a generic configuration from a MD trajectory at $T=300$ K.
@@
All planar capacitor samples are enclosed in an orthorhombic box, with periodic boundary conditions applied. More precisely, at first samples are periodically replicated in 2D along the $(yz)$ plane. Then, despite the inherent 2D character of the capacitor, the electrostatic interactions have been computed using 3D Ewald sums upon replicating the capacitor also along the $\hat{x}$ direction, because 2D Ewald sum routines are generally considered to be less efficient than their 3D counterparts. No explicit test, however, has been carried out to quantify or even verify this statement for the present case. the periodicity along $x$ is equal to the combined width of the graphite (9-layer) block and of the [bmim][NTf2] slab. In other words, no gap is left between the sample and its nearest replicas along $x$. The thickness of the graphite slab ($\sim 30$ Å ) is the major aspect that limits the spurious interactions of the charged interfaces and [bmim][NTf2] slab across the pbc boundary.
The artificial periodicity along $x$, however, has a far more significant effect, that needs to be accounted for. An isolated, single planar capacitor of charge $Q=\sigma A$, where $\sigma$ is the surface charge density on the positive electrode and $A$ is the electrode area, is globally neutral. Hence, the electric field on its outside vanishes, and symmetry considerations show that inside the capacitor the electric field is oriented normal to the plates. Then, application of Gauss’ theorem immediately gives that the electric field strength on its inside is a constant equal to $4\pi\sigma$. For an infinitely periodic sequence of capacitors, however, the analysis is not so simple, because there is no simple way to locate a surface of vanishing electric field (${\bf E}=0$) from which to start applying Gauss’ theorem to compute ${\bf E}(x)$ anywhere else in space. Seen in a different but equivalent way, there is no reason why the electric field within the graphite blocks separating neighbouring replicas has to vanish. For these reasons, the electric field inside the liquid electrolyte, although still constant and directed along $x$, is not equal to $4\pi\sigma$.chris The precise value of ${\bf E}(x)$ can be computed analytically based on the Ewald sum formalism, but it depends on boundary condition parameters set in the simulation code, whose value in not easily determined. To avoid this ambiguity, the electric field $^{\bf E_{0}}=E_{0}(x)\hat{x}$ along the $x$ coordinate within the capacitor and within the graphite slab has been computed using Gromacs considering an empty capacitor whose charged plates (represented by the charge graphite layers) are located at the distance used in the actual simulations with the ionic liquid. This empty capacitor curve $E_{0}(x)$ has been used throughout the paper to compute the electrode contribution to the position-dependent electric field ${\bf E}(x)$ and electrostatic potential ${\bf\phi}(x)$ through the capacitor.
@@
All simulated systems, including the homogeneous liquid [bmim][NTf2] samples, consist of the same number of cations and anions. The electrolyte slab, therefore, is globally neutral. Since also the whole sample, which is 3D periodic, needs to be neutral, the two interfaces delimiting the electrolyte may only carry zero or opposite surface charge densities $\pm\sigma$. At first, the charge on each graphite C atom is identically zero. Then, any non-vanishing charge density on the graphite surfaces will be represented by atomic charges $q_{grf}=\pm\delta q$ of equal modulus and opposite sign on the two surfaces. Since graphite is a conductor, the static $\pm\sigma$ charges on the two sides of the graphite, in reality, would not be stable with respect to charge recombination. However, the electrical conductivity perpendicular to the graphite planes is about $10^{3}$ times lower than along the planes. Therefore, the simple charging model is justified by neglecting the electrical conductivity across the graphite slab, and assuming that the two outermost C-planes are connected to the two poles of a battery.
All simulations have been carried out using the Gromacsgromacs simulation package, version 2019.3. The analysis of trajectories has been made using home-made programs. These have been used to compute number and charge density profiles along the direction normal to the graphite planes, ion-ion radial distribution functions, self-diffusion coefficient based on the Einstein equation, orientation correlations among the ions and with respect to the planar IL/graphite interfaces. The average electrostatic potential along the same normal direction has been computed by integrating the electric field inside he capacitor, which is the sum of the [bmim][NTf2] contribution determined from the average charge density profile of the ions, and the contribution from the charge on the electrodes, computed by Gromacs for the empty capacitor, as explained a few paragraphs above in this same section.
Figure 4: Snapshot of the ionic liquid part of Fig. 1 coarse grained through Eq. 1. Red dots: cations; blue dots: anions. Temperature in measured in scaled energy units, see text.
@@
whose definition is apparently analogous to that of the centre of mass of a set of particles. The choice of emphasising the charges instead of the masses is due to our interest in the electrostatics of the sample. This definition is not as unanimously accepted nor extensively used as that of the centre of mass. In this context, however, it provides a simple way to coarse grain the charge distribution of [bmim][NTf2], which is fairly complex and difficult to visualise and interpret in its atomistic detail. For instance, the configuration of cations and anions corresponding to the snapshot in Fig. 1 is shown in Fig. 4, highlighting the drastic simplification of the IL structure achieved by the coarse graining expressed by Eq. 1.
@@ III.1 Properties of the homogeneous samples
Since the experimental density of [bmim][NTf2] at T=300T=300K was fitted by increasing the LJ radii, the computed average density of the model, $\rho=1.4602\pm 0.0002$ g/cm3, falls into the range of experimental values (from $\rho=1.44$ to $1.47$ g/cm3) measured in experiments at the same temperature. The thermal expansion coefficient is somewhat underestimated, being $K_{T}=2.72\ 10^{-4}$ K-1 for the model, versus the experimental $K_{T}=6.4\ 10^{-4}$ K-1. At the $20$ molar % concentration, water appears to be soluble in the re-parametrised [bmim][NTf2] model, in agreement with experiments.
Figure 5: Number ($\rho_{N}(x)$, blue line) and charge ($\rho_{Q}(x)$, red line) density profiles across the uncharged ($\sigma=0$) capacitor. $L_{x}$ is the average separation of the outermost graphite layers measured across the [bmim][NTf2] slab.
@@
Figure 6: Number ($\rho_{N}(x)$, blue line) and charge ($\rho_{Q}(x)$, red line) density profiles across the capacitor with $\sigma=2$ e/nm2 surface charge density. $L_{x}$ is the average separation of the outermost graphite layers measured across the [bmim][NTf2] slab. Notice the change of scale with respect to Fig. 5.
@@ III.2 Structural, dynamic and electrostatic properties of the graphite / [bmim][NTf2] interface
The reference frame has been oriented in such a way that the interface is orthogonal to the $x$ axis. For this reason, local properties such as density and electrostatic potential profiles are plotted as a function of $x$. To simulate charged capacitors, small charges of up to a few $10^{-2}$ e (where e is the H+ charge) have been equally assigned to all the atoms of the two most external layers of graphite, qualitatively reflecting the picture provided by ab-initio (density functional theory) computations, see charging, . The charges are of equal magnitude and opposite sign on the two charged layers. Since also the IL slab is globally neutral, the whole sample is neutral. Since, according to the simulation results, the 2D number density of C atoms in the graphite layers at $T=300$ K is $37.08$ nm-2, a charge density of 0.05 e per graphite layer corresponds to a sizeable surface charge density of nearly $2$ e/nm-2.
Figure 7: Electric field and electrostatic potential as a function of the $x$ coordinate across the capacitor for two different values of the surface charge $\sigma$. The gray rectangles represent the graphite electrodes. $L_{x}$ is the average separation of the outermost graphite layers measured across the [bmim][NTf2] slab. The scaled units for $E(x)$ and $\varphi(x)$ are explained in the text.
@@ III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
The results of the previous subsections show that the average electric field vanishes away from the interfaces. The presence of electro-optical effects sufficiently strong to be detected by spectroscopy, however, does not require a non-vanishing static electric field. A detectable Stark effect, for instance, can result from a change in the average square modulus $E^{2}$ of the fluctuating electric field acting on each ion. Such a square modulus does not vanish even in bulk [bmim][NTf2], and changes in its value can result from subtle correlations compatible with the strict screening sum rules that exclude net charge separation and non-vanishing average electric field away from the interface. Thus, excluding the violation of these rules, the present investigation focuses on possible correlations among electrostatic dipoles.
@@
Such an analysis can follow a variety of leads, but two aspects need to be retained. First, one needs to define dipoles inside an ionic fluid, which is certainly a challenging undertaking since the [bmim]+ and [NTf2]- ions are mobile. Second, each dipole needs to have a compact geometry in such a way that its position in space can be specified with good precision, allowing the meaningful computation of a local density of dipoles everywhere through the system as well as the definition of distance dependent correlation functions. The subdivision of the system charges into dipoles cannot, but also does not need to be permanent. In fact, the identity of the ions forming any given dipole will evolve with the changing configuration of ions in the ionic liquid. However, to lend credibility to the physical relevance of these dipoles, their lifetime should be long. Moreover, the probability distribution for the modulus $R_{d}$ over the dipoles population should be relatively narrow, and correspond to a charge separation comparable to the average distance of ions as given by a ion-ion pair distribution function.
@@
With these considerations in mind, we set out to partition the moving charges of the system into a population of dipoles. The dipole moment of a set of charges is unambiguously defined provided the aggregation of charge is overall neutral. Therefore, we focus our attention on the dipole of neutral ion pairs, consisting of the association of one cation and one anion. The difficulty with this approach is that the way to partition an assembly of ions into a collection of dipoles by joining a [bmim]+ cation and a neighbouring [NTf2]- anion is not unique. The most intuitive approach could be to join each ion with its nearest neighbour of the opposite sign. However, the property of ion A+ being the nearest neighbour of B- is not symmetric, since the nearest cation of B- might be C+, different from A+. Simple algorithms like starting from a given ion, joining it to the closest counter-ion not already engaged into a dipole do not solve the problem. At first, this approach provides a series of tightly bound dipoles, satisfying the requirements listed above. Soon, however, going down the series of cations, the inherent frustration in the process of pairing neighbouring ions becomes manifest since the algorithm increasingly joins pairs separated by long distances, making meaningless the location of the dipole in space. Moreover, the result would depend on the ordering in which the ions are paired, since the first pairs to be formed are systematically favoured (i.e., have shorter charge separation) with respect to the last.
@@
The approach that is proposed and tested here is to define such a unique partition of the ion population into dipoles by requesting that the total self-energy of all dipoles is a minimum among all possible subdivision of all cations and anions into neutral pairs.
Figure 8: Snapshots of the dipole configurations at the two extreme temperatures of the simulated annealing (see text) of the dry capacitor at $\sigma=0$. The $x$ axis is orthogonal to the two IL - graphite interfaces, with graphite not shown in the picture. Dark gray arrows point to the left, red arrows point to the right. $T$ is a fictitious temperature introduced to drive the MC optimisation of the ions’ pairing into dipoles, not to be confused with the physical temperature (T=300 K) of the graphite - liquid [bmim][NTf2] simulation. The units of the fictitious temperature are discussed in the text.
Let us first consider the case of neutral graphite surfaces. The case of charged interfaces will require some slight modifications. The self-energy of a dipole made of charges (q,q)(q,-q) separated by a distance $s$ usually is given as $u_{s}=-q^{2}/s$. In the present case, in which cations and anions consists of several atoms joined into a complex ion, the self-energy of the dipole made of cation $\alpha$ and anion $\beta$ is defined as:
@@
where $i$ and $j$ label atoms (of charges $q_{i}$ and $q_{j}$, respectively) belonging to cation $\alpha$ and anion $\beta$, respectively. A very similar (although not identical) value of the self-energy is obtained by approximating $u_{s}(\alpha,\beta)$ with:
$\tilde{u}_{s}(\alpha,\beta)=-\frac{1}{\mid{\bf R}^{(Q)}_{\alpha}-{\bf R}^{(Q)}_{\beta}\mid}$ (4)
where ${\bf R^{(Q)}}_{\alpha}$, ${\bf R^{(Q)}}_{\beta}$ are defined in Eq. 1. The good correspondence between $u_{s}(\alpha,\beta)$ and $\tilde{u}_{s}(\alpha,\beta)$ confirms the ability of Eq. 1 to provide a meaningful coarse grained description of the charge distribution in the [bmim][NTf2] slab, and could ease the analysis of configurations selected from the simulation of large capacitor.
On the basis of these considerations and observations, a procedure has been devised to carry out the unbiased subdivision of any given configuration of ions into a configuration of dipoles, consisting of:
  • start from the pairing of each cation with its nearest anion which is not already engaged in another pair;